total intensity of earth's magnetic field at magnetic equator is 5 units what is the value at  a magnetic latitude of 37 degree

Grade:12

1 Answers

Nikhil
4 years ago
As given let mag field at equator is 5 units  ,
So mag field at poles is 10
Now as the component of vertical is cos37
So vertical is 10×4/5 =8 units
And horizontal is 5×3/5=3 units
So total B=(8^2+3^2)^1/2=(73)^1/2